Please solve and show all work. It would be greatly appreciated if you explained it too. Thanks.
Divide:
4x^2 - 15 / 2x - 5

Found 2 solutions by stanbon, Theo:
Answer by stanbon(75887)

Please solve and show all work. It would be greatly appreciated if you explained it too. Thanks.
Divide:
4x^2 - 15 / 2x - 5
----
Set it up like a long-division problem
---
Divide the 4x^2 by 2x to get 2x (that is the 1st term of the quotient)
---
Multiply 2x(2x-5) = 4x^2-10x
----
Subtract that from 4x^2-15 to get: 10x-15
----
Divide that by 2x to get 5 (that is the 2nd term of the quotient
-----------
Multiply 5*(2x-5) = get 10x-25
---
Subtract that from 10x-15 to get: 10 (that is the remainder)
Answer:
Quotient: 2x+5
Remainder: 10
=======================
cheers,
Stan H.
=========================

Answer by Theo(13342)

you want to divide 4x^2 - 15 by (2x - 5)
you need to fill in the missing orders of exponents.
when you do that, your equation you are dividing into will be equal to:
4x^2 + 0x - 15
you want to divide that by (2x - 5)
you start by dividing 2x into 4x^2 to get 2x because 2x * 2x = 4x^2
you then multiply (2x - 5) by 2x to get 4x^2 - 10x
you then subtract 4x^2 - 10x from 4x^2 + 0x - 15 to get 10x - 15
you then divide 2x into 10x to get 5 because 5 * 2x = 10x
you then multiply (2x - 5) by 5 to get 10x - 25
you then subtract 10x - 25 from 10x - 15 to get + 10
since + 10 is of a lower order than 2x, you are done and your remainder is 10.
your quotient is 2x + 5 + 10/(2x - 5)

instructions on how to do polynomial division can be found at the following website.
http://www.wtamu.edu/academic/anns/mps/math/mathlab/col_algebra/col_alg_tut36_longdiv.htm

instructions on how to do synthetic division can be found at the following website.
http://www.wtamu.edu/academic/anns/mps/math/mathlab/col_algebra/col_alg_tut37_syndiv.htm

email: theoptsadc@yahoo.com
website: http://theo.x10hosting.com